rabby1:

Sydney bought a 4-pound bag of brown rice at the grocery store. The nutritional information showed that the bag contained 37 servings. If there are 16 ounces in one pound, exactly how many ounces are there in one serving of brown rice?

jhonyy9:

use the rule of 3 simple and so you can writing 1 pound contain 16 ounces 4 pound contain x ounces now than you cross multiplie will get : 1*x = 4*16 so x = ? now you know that these 4 pound bag contain x quantity ounces what contained 37 servings so now you can using the same like in the first step the rule of 3 simple and write 37 services contain x quantity ounces 1 service contain y quantity ounce cross multiply and get y = 1*x/37 so this will be the quantity ounce what contain one service hope helped understandably